December 7, 2020, Revision: 1.92

1

            Figure 1: Isothermal

$Header: /var/cvs/brandenb/tex/etc/notes.tex,v 1.92 2020/11/22 12:17:57 brandenb Exp $

                                                    Figure 2: Isothermal_cs_tdep